# Classical and quantum thermodynamics in a non–equilibrium regime: Application to Stirling engine

### v1

@@ Appendix B DL non–equilibrium minimum work principle
Figure 6: Schematic to derive Eq. (51). The blue circles 1 and 2 represent the equilibrium states, and the red square A represents the non–equilibrium state. The quasi-static work performed for $1\rightarrow 2$ (red dashed line) is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow 2}^{\rm min}$, while the non–equilibrium work from 1 to A (dark blue curve) is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$. When $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$ is fixed, the value of $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ depends on the path (green and light blue curves). From the second law of thermodynamics, we have $(W_{\rm A}^{ext})_{1\rightarrow A}+(\tilde{W}_{\rm A}^{ext})_{\rm A\rightarrow 2}\geq-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}$. This indicates a lower bound of work for ${A\rightarrow 2}$ denoted by $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ (light blue curve) and given by Eq. (51).
@@
Here, we derive fundamental equations to develop the DL non–equilibrium minimum work principle. Consider two equilibrium states 1 and 2 that are connected by a quasi-static process (see Fig. 6). From the principle of DL minimum work [Eq. (8)], the work done in this process is equivalent to the difference in Massieu potentials: $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow 2}^{\rm min}=-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}\equiv(\Phi_{\rm A}^{\rm qst})_{1}-(\Phi_{\rm A}^{\rm qst})_{2}$, where ${n\rightarrow n^{\prime}}$ represents the transition from any state $n$ to $n^{\prime}$. Separately, we consider the non–equilibrium state $A$ and introduce the non–equilibrium process $1\rightarrow A$ (dark blue curve in Fig. 6), and any process from $A$ to 2 (green and light blue curves). For each of these processes, the work done is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$ and $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$, respectively. For fixed $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$, the value of $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ changes depending on the path.
@@
From the second law of thermodynamics [Eq. (8)], we then have $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}+(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}>-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}.$ This indicates that $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ has a lower bound for fixed $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$; otherwise the inequality is violated, for example, as $-\infty>-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}-(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$. Using the lower bound of the work, we introduce the non–equilibrium Massieu potential $\Phi_{\rm A}^{\rm neq}$ defined as
@@
Figure 7: Schematic to derive the principle of DL non–equilibrium minimum work expressed by Eq. (53). We consider two non–equilibrium states $A$ and $B$ (red square) that are connected by a non–equilibrium process described by the red arrow for $A\rightarrow B$. The light blue curves $A\rightarrow 2$ and $B\rightarrow 2$ represent non–equilibrium-to-equilibrium minimum work paths whose work is denoted by $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ and $(\tilde{W}_{\rm A}^{ext})_{B\rightarrow 2}^{\rm min}$, respectively.
Introducing a second non–equilibrium state $B$ on the pathway $A\rightarrow 2$ as depicted in Fig. 7, we now discuss the non–equilibrium transition ${A\rightarrow B}$ (red arrow). We consider the non–equilibrium-to-equilibrium minimum work from $B$ to 2 (light blue curve), expressed as
@@
From the inequality $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}\geq-[(\Phi_{\rm A}^{\rm qst})_{2}-(\Phi_{\rm A}^{\rm neq})_{A}]$ and Eq. (52), we have
@@
Because a non–equilibrium process changes as a function of time, it is more convenient to use the time-derivative form of the above. Thus, for the process from $A$ to $B$ at times $t$ and $t+dt$, where $dt$ is an infinitesimal time, we have the following inequality:
@@
This inequality is an extension of the principle of minimum work to a non–equilibrium regime.
Figure 8: Schematic showing that the non–equilibrium Massieu potential can be evaluated from any equilibrium state 3 instead of 2. To illustrate this, we consider the non–equilibrium-to-equilibrium path $A\rightarrow 3$, where we have introduced a third equilibrium state denoted by 3 (blue circle). The work between the two equilibrium states is denoted by $(W_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}$ and $(W_{\rm A}^{ext})_{3\rightarrow 2}^{\rm qst}$, respectively.
Although we have derived Eqs. (53) and (54) for the specific equilibrium state 2, actually the value $(\Phi_{\rm A}^{\rm neq})_{A}$ is the same for any quasi-equilibrium state along $\beta^{\rm qst}(t)$ and $x^{\rm qst}(t)$. To illustrate this, we introduce a third equilibrium state 3 and consider the transition $A\rightarrow 3$ (see Fig. 8). Because $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}+(\tilde{W}_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}\geq(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}$ for $A\rightarrow 2\rightarrow 3$ and $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}+(\tilde{W}_{\rm A}^{ext})_{3\rightarrow 2}^{\rm qst}\geq(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ for $A\rightarrow 2\rightarrow 3$, we have
@@
where we have used $(W_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}=(\Phi_{\rm A}^{\rm qst})_{3}-(\Phi_{\rm A}^{\rm qst})_{2}$. From Eq. (51), this indicates that the non–equilibrium Massieu potential is independent of the choice of the equilibrium state 2.
Because the other potentials are interrelated through TDL transformations, we have
@@
However, unlike in the Massieu–Planck case, the equality in Eqs. (59) and (60) cannot be obtained in general because the minimum DL work path in Fig. 6 is evaulaed by optimizing both $\beta(t)$ and $x(t)$, while $\beta$ has been fixed to introduce the Helmholtz energy as $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta$.
The above discussion indicates that by including the thermostatic process, we may develop an efficient heat engine under non–equilibrium conditions.
Appendix C Convexity of non–equilibrium thermodynamic potentials as functions of extensive variables
Table 7: Inequalities holding between the non–equilibrium state at time $t_{1}$ and the equilibrium state at time $t_{2}$, together with the conditions under which these inequalities hold. From these inequalities, we find that, for example, for a given inverse temperature and electric field, the non–equilibrium Planck potential takes its maximum value when the state is in equilibrium.
@@ Appendix C Convexity of non–equilibrium thermodynamic potentials as functions of extensive variables
Figure 9: Schematic for deriving convexity relations. The red squares represent the non–equilibrium states. The arrows are the relaxation paths to equilibrium state 2 described by the time-evolution operator $\hat{U}_{\rm A}(t,t_{0})$: The blue arrow is the DL minimum work path from state $\lambda$.
To obtain stable thermodynamic properties, the thermodynamic potential must be a convex function of the work variables. Even in the non–equilibrium case, we find that convexity holds under certain conditions. Here, we show this in the case of the SB model. Because other potentials can be discussed similarly, we limit our discussion here to the case of a non–equilibrium Massieu potential.
We consider three non–equilibrium states $A$, $B$, and $\lambda$ ($\lambda\in[0,1]$) with the same initial inverse temperature $\beta(t_{0})$ at time $t_{0}$. We then assume that the state $\lambda$ satisfies the relation
@@
$(W_{\rm A}^{ext})_{\lambda}(t,t_{0})=(1-\lambda)(W_{\rm A}^{ext})_{A}(t,t_{0})+\lambda(W_{\rm A}^{ext})_{B}(t,t_{0}).$ (64)
The DL minimum work path from the state $\lambda$ to the equilibrium state 2 is described by $\hat{U}_{\rm A}(t,t_{0})$ (see Fig. 9). This indicates that when $t$ is large, the intensive variables must be $\beta(t)\rightarrow\beta_{2}$ and ${x}(t)\rightarrow x_{2}$, where $\beta_{2}$ and $x_{2}$ are the values at the equilibrium state 2. Therefore, the processes starting from $A$ and $B$ also relax to the equilibrium state 2 as $t\rightarrow\infty$. Then, from Eq. (64), we obtain
@@
Using the inequality $(W_{\rm A}^{ext})_{\alpha\rightarrow 2}\geq(W_{\rm A}^{ext})_{\alpha\rightarrow 2}^{\rm min}$ for the states $\alpha=A$ and $B$, and Eq. (51), we obtain the inequality to prove convexity as
@@
where $(\Phi_{\rm A}^{\rm neq})_{\alpha}$ is the non–equilibrium Massieu potential in the state $\alpha$ ($\alpha=A$, $B$, and $\lambda$). When $A$ and $B$ are in equilibrium states, we obtain the convexity relation in the quasi-static case for fixed inverse temperature using the inequality for the Massieu potential in Table 8 as Koyanagi and Tanimura (2024a)
@@
We summarize the convexity and concavity properties for the other non–equilibrium thermodynamic potential in Table 7.
Appendix D Inequality between non–equilibrium and quasi-static thermodynamic potentials
Table 8: Inequalities for Massieu–Planck potentials holding between the non–equilibrium state $A$ at time $t_{1}$ and the equilibrium state 2 at time $t_{2}$, together with the conditions under which these inequalities hold. From these inequalities, we find that, for example, for a given inverse temperature and intensive variable, the non–equilibrium Planck potential takes its maximum value when $A$ is in the equilibrium state.
@@ Appendix D Inequality between non–equilibrium and quasi-static thermodynamic potentials
Table 9: Inequalities for Helmholtz–Gibbs potentials holding between the non–equilibrium state at time $t_{1}$ and the equilibrium state at time $t_{2}$, together with the conditions under which these inequalities hold.
@@
Under given conditions, the non–equilibrium DL thermodynamic potentials are smaller than the quasi-static ones. Here, we derive the inequalities between the quasi-static and non–equilibrium thermodynamic potentials.
We consider a non–equilibrium-to-equilibrium transition $A\rightarrow 2$ from time $t_{1}$ to $t_{2}$ described by the inverse temperature $\beta(t)$ and the extensive variable $X_{\rm A}^{\rm neq}(t)$. The DL extensive work in this process is evaluated as
@@
where we have assumed that $\beta(t_{1})=\beta_{2}$ and $X_{\rm A}^{\rm neq}(t_{1})=(X_{\rm A}^{\rm qst})_{2}$, and that $\beta_{2}=\beta(t)$ and $\tilde{x}_{2}=\tilde{x}(t)$ for time $t>t_{1}$, to make the subsystem relax to the equilibrium state 2 at time $t_{2}$. Because $X_{\rm A}^{\rm neq}(t_{2})=(X_{\rm A}^{\rm qst})_{2}=X_{\rm A}^{\rm neq}(t_{1})$, the right-hand side of Eq. (D) vanishes. Thus, the net DL extensive work is zero, and so the DL minimum work principle reduces to $0\geq(\Phi_{\rm A}^{\rm neq})_{A\rightarrow 2}$, from which it follows that $\Phi_{\rm A}^{\rm qst}(t_{2})\geq\Phi_{\rm A}^{\rm neq}(t_{1})$.
The inequalities for the other DL thermodynamic potentials are obtained in the same manner. We summarize the results in Tables 8 and 9
Figure 10: Difference between the quasi-static and non–equilibrium Planck potential in the quantum and classical cases (solid and dashed curves, respectively) for one cycle of the Stirling engine. The blue, green, and red curves represent the cases of weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling, respectively.
In Fig. 10, to illustrate the inequalities presented in Table 8, we plot the difference between the quasi-static and non–equilibrium Planck potential, $\Delta\Xi_{\rm A}(t)=\Xi_{\rm A}^{\rm qst}[\beta(t),\tilde{x}(t)]-\Xi_{\rm A}^{\rm neq}(t)$, for one cycle of the Stirling engine described in Sec. III.4. As can be seen, $\Delta\Xi_{\rm A}(t)$ is always positive and satisfies the inequalities in Table 8.

### v2

@@ Appendix B DL non-equilibrium minimum work principle
Figure 6: Schematic to derive Eq. (51). The blue circles 1 and 2 represent the equilibrium states, and the red square A represents the non-equilibrium state. The quasi-static work performed for $1\rightarrow 2$ (red dashed line) is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow 2}^{\rm min}$, while the non-equilibrium work from 1 to A (dark blue curve) is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$. When $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$ is fixed, the value of $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ depends on the path (green and light blue curves). From the second law of thermodynamics, we have $(W_{\rm A}^{ext})_{1\rightarrow A}+(\tilde{W}_{\rm A}^{ext})_{\rm A\rightarrow 2}\geq-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}$. This indicates a lower bound of work for ${A\rightarrow 2}$ denoted by $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ (light blue curve) and given by Eq. (51).
@@
Here, we derive fundamental equations to develop the DL non-equilibrium minimum work principle. Consider two equilibrium states 1 and 2, which are connected by a quasi-static process (see Fig. 6). From the principle of DL minimum work [Eq. (8)], the work performed in this process is equivalent to the difference in the Massieu potentials: $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow 2}^{\rm min}=-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}\equiv(\Phi_{\rm A}^{\rm qst})_{1}-(\Phi_{\rm A}^{\rm qst})_{2}$, where ${n\rightarrow n^{\prime}}$ represents the transition from any state $n$ to $n^{\prime}$. Separately, we consider the non-equilibrium state $A$ and introduce the non-equilibrium process $1\rightarrow A$ (dark blue curve in Fig. 6), and any process from $A$ to 2 (green and light blue curves). For each of these processes, the work performed is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$ and $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$, respectively. For fixed $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$, the value of $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ changes depending on path.
@@
From the second law of thermodynamics [Eq. (8)], we obtain $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}+(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}>-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}.$ This indicates that $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ has a lower bound for fixed $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$; otherwise, the inequality is violated, for example, as follows: $-\infty>-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}-(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$. Using the lower bound of this work, we introduce the non-equilibrium Massieu potential $\Phi_{\rm A}^{\rm neq}$ defined as
@@
Figure 7: Schematic for deriving the principle of DL non-equilibrium minimum work expressed by Eq. (53). We consider two non-equilibrium states $A$ and $B$ (red square) that are connected by a non-equilibrium process described by the red arrow for $A\rightarrow B$. The light blue curves $A\rightarrow 2$ and $B\rightarrow 2$ represent non-equilibrium-to-equilibrium minimum work paths, whose work is denoted by $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ and $(\tilde{W}_{\rm A}^{ext})_{B\rightarrow 2}^{\rm min}$, respectively.
By introducing a second non-equilibrium state $B$ on the pathway $A\rightarrow 2$ as depicted in Fig. 7, we now discuss the non-equilibrium transition ${A\rightarrow B}$ (red arrow). We now consider the non-equilibrium-to-equilibrium minimum work from $B$ to 2 (light blue curve), expressed as
@@
From the inequality $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}\geq-[(\Phi_{\rm A}^{\rm qst})_{2}-(\Phi_{\rm A}^{\rm neq})_{A}]$ and Eq. (52), we obtain
@@
Because a non-equilibrium process changes as a function of time, it is more convenient to use the time derivative form of the above. Thus, for the process from $A$ to $B$ at times $t$ and $t+dt$, where $dt$ is infinitesimal time, we obtain the following inequality:
@@
This inequality extends the principle of minimum work to a non-equilibrium regime.
Figure 8: Schematic showing that the non-equilibrium Massieu potential can be evaluated from any equilibrium state 3 instead of 2. To illustrate this, we consider the non-equilibrium-to-equilibrium path $A\rightarrow 3$, where we introduce a third equilibrium state denoted by 3 (blue circle). The work between the two equilibrium states is denoted by $(W_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}$ and $(W_{\rm A}^{ext})_{3\rightarrow 2}^{\rm qst}$, respectively.
Although we have derived Eqs. (53) and (54) for the specific equilibrium state 2, the value $(\Phi_{\rm A}^{\rm neq})_{A}$ is the same for any quasi-equilibrium state along $\beta^{\rm qst}(t)$ and $x^{\rm qst}(t)$. To illustrate this, we introduce a third equilibrium state 3 and consider the transition $A\rightarrow 3$ (see Fig. 8). Because $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}+(\tilde{W}_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}\geq(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}$ for $A\rightarrow 2\rightarrow 3$ and $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}+(\tilde{W}_{\rm A}^{ext})_{3\rightarrow 2}^{\rm qst}\geq(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ for $A\rightarrow 3\rightarrow 2$, we have
@@
where we have used $(W_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}=(\Phi_{\rm A}^{\rm qst})_{3}-(\Phi_{\rm A}^{\rm qst})_{2}$. From Eq. (51), this indicates that the non-equilibrium Massieu potential is independent of the choice of the equilibrium state 2.
Because the other potentials are interrelated through TDL transformations, we obtain
@@
However, unlike in the Massieu–Planck case, the equality in Eqs. (59) and (60) cannot be obtained in general because the minimum DL work path in Fig. 6 is evaluated by optimizing both $\beta(t)$ and $x(t)$, while $\beta$ has been fixed to introduce the Helmholtz energy as $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta$.
The above discussion indicates that the inclusion of the thermostatic process is a key to develop an efficient heat engine under non-equilibrium conditions.
Appendix C Convexity of non-equilibrium thermodynamic potentials as functions of extensive variables
Table 7: Inequalities between the non-equilibrium state at time $t_{1}$ and the equilibrium state at time $t_{2}$, with the conditions under which these inequalities hold. From these inequalities, we find that, for example, for a given inverse temperature and external field, the non-equilibrium Planck potential takes its maximum value when the state is in equilibrium.
@@ Appendix C Convexity of non-equilibrium thermodynamic potentials as functions of extensive variables
Figure 9: Schematic for deriving convexity relations. The red squares represent the non-equilibrium states. The arrows are the relaxation paths to equilibrium state 2, as described by the time-evolution operator $\hat{U}_{\rm A}(t,t_{0})$: The blue arrow is the DL minimum work path from state $\lambda$.
To obtain stable thermodynamic properties, the thermodynamic potential must be a convex function of the work variables. Even in the non-equilibrium case, we find that convexity holds under certain conditions. Here, we present this in the SB model. Because other potentials can be discussed similarly, we have limited our discussion to the case of non-equilibrium Massieu potential.
We consider three non-equilibrium states $A$, $B$, and $\lambda$ ($\lambda\in[0,1]$) with the same initial inverse temperature $\beta(t_{0})$ at time $t_{0}$. We then assume that the state $\lambda$ satisfies the relation
@@
$(W_{\rm A}^{ext})_{\lambda}(t,t_{0})=(1-\lambda)(W_{\rm A}^{ext})_{A}(t,t_{0})$
$+\lambda(W_{\rm A}^{ext})_{B}(t,t_{0}).$ (64)
The DL minimum work path from the state $\lambda$ to the equilibrium state 2 is described by $\hat{U}_{\rm A}(t,t_{0})$ (see Fig. 9). This indicates that when $t$ is large, the intensive variables must be $\beta(t)\rightarrow\beta_{2}$ and ${x}(t)\rightarrow x_{2}$, where $\beta_{2}$ and $x_{2}$ are the values at the equilibrium state 2. Therefore, processes starting from $A$ and $B$ also relax to the equilibrium state 2 as $t\rightarrow\infty$. Then, from Eq. (C), we obtain
@@
Using the inequality $(W_{\rm A}^{ext})_{\alpha\rightarrow 2}\geq(W_{\rm A}^{ext})_{\alpha\rightarrow 2}^{\rm min}$ for the states $\alpha=A$ and $B$ and Eq. (51), we obtain the inequality to prove convexity as
@@
where $(\Phi_{\rm A}^{\rm neq})_{\alpha}$ is the non-equilibrium Massieu potential in the state $\alpha$ ($\alpha=A$, $B$, and $\lambda$). When $A$ and $B$ are in equilibrium, we obtain the convexity relation in the quasi-static case for a fixed inverse temperature using the inequality for the Massieu potential as Table 8 as Koyanagi and Tanimura (2024b)
@@
The convexity and concavity properties for the other non-equilibrium thermodynamic potentials are summarized in Table 7.
Appendix D Inequality between non-equilibrium and quasi-static thermodynamic potentials
Table 8: Inequalities for Massieu–Planck potentials between the non-equilibrium state $A$ at time $t_{1}$ and the equilibrium state 2 at time $t_{2}$, together with the conditions under which these inequalities hold. From these inequalities, we find that, for example, for a given inverse temperature and intensive variable, the non-equilibrium Planck potential takes its maximum value when $A$ is in the equilibrium state.
@@ Appendix D Inequality between non-equilibrium and quasi-static thermodynamic potentials
Table 9: Inequalities for Helmholtz–Gibbs potentials between the non-equilibrium state at time $t_{1}$ and the equilibrium state at time $t_{2}$, with the conditions under which these inequalities hold.
@@
Under given conditions, the non-equilibrium DL thermodynamic potentials are smaller than the quasi-static ones. Here, we derive the inequalities between the quasi-static and non-equilibrium thermodynamic potentials.
We consider a non-equilibrium-to-equilibrium transition $A\rightarrow 2$ from time $t_{1}$ to $t_{2}$ described by the inverse temperature $\beta(t)$ and the extensive variable $X_{\rm A}^{\rm neq}(t)$. The DL extensive work in this process is evaluated as
@@
where we have assumed that $\beta(t_{1})=\beta_{2}$ and $X_{\rm A}^{\rm neq}(t_{1})=(X_{\rm A}^{\rm qst})_{2}$, and that $\beta_{2}=\beta(t)$ and $\tilde{x}_{2}=\tilde{x}(t)$ for time $t>t_{1}$, allowing the subsystem to relax to the equilibrium state 2 at time $t_{2}$. Because $X_{\rm A}^{\rm neq}(t_{2})=(X_{\rm A}^{\rm qst})_{2}=X_{\rm A}^{\rm neq}(t_{1})$, the right-hand side of Eq. (D) vanishes. Thus, the net DL extensive work is zero, and so the DL minimum work principle reduces to $0\geq(\Phi_{\rm A}^{\rm neq})_{A\rightarrow 2}$, from which it follows that $\Phi_{\rm A}^{\rm qst}(t_{2})\geq\Phi_{\rm A}^{\rm neq}(t_{1})$.
The inequalities for the other DL thermodynamic potentials were obtained in the same manner. The results are summarized in Tables 8 and 9
Figure 10: Difference between the quasi-static and non-equilibrium Planck potentials in the quantum and classical cases (solid and dashed curves, respectively) for one cycle of the thermostatic Stirling engine. The blue, green, and red curves represent the weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling, respectively.
In Fig. 10, to illustrate the inequalities presented in Table 8, we plot the difference between the quasi-static and non-equilibrium Planck potential, $\Delta\Xi_{\rm A}(t)=\Xi_{\rm A}^{\rm qst}[\beta(t),\tilde{x}(t)]-\Xi_{\rm A}^{\rm neq}(t)$, for one cycle of the thermostatic Stirling engine described in Sec. III.4. As can be seen, $\Delta\Xi_{\rm A}(t)$ is always positive and satisfies the inequalities in Table 8.